The horn of Ms Brown’s VW has a frequency of 400 Hz. What frequency would her 3U math class hear if she honks the horn while driving at 50 km/h on a 23 C day?

a) toward the class

b) away from the class


1
Expert's answer
2021-06-22T13:33:27-0400

a) "\nu=\nu_0\frac{c}{c-u}=400\cdot\frac{345.3}{345.3-13.89}=416.76\ (Hz)"


b) "\nu=\nu_0\frac{c}{c+u}=400\cdot\frac{345.3}{345.3+13.89}=384.53\ (Hz)"